I am completely lost on how to work this problem. solve:
y-2-8sqrt(y-2)+15 = 0

Answer by scott8148(5891)

get the sqrt term by itself and then square both sides ... y+13=8sqrt(y-2) ... (y+13)^2=64(y-2)

y^2+26y+169=64y-128 ... y^2-38y+297=0 ... factoring gives (y-27)(y-11)=0 ... y=27, y=11